# CodeR: Issue Resolving with Multi-Agent and Task Graphs

### v1

@@ Abstract
GitHub issue resolving recently has attracted significant attention from academia and industry. SWE-bench [1] is proposed to measure the performance in resolving issues. In this paper, we propose CodeR, which adopts a multi-agent framework and pre-defined task graphs to Repair & Resolve reported bugs and add new features within code Repository. On SWE-bench lite, CodeR is able to solve $28.00\%$ of issues, in the case of submitting only once for each issue. We examine the performance impact of each design of CodeR and offer insights to advance this research direction [^1] https://github.com/NL2Code/CodeR.
@@ 1 Introduction
The rapidly growing capability of Large Language Models (LLMs) is dramatically reshaping many industries [2, 3, 4]. The most recent release of GPT-4o [5] demonstrates a significant leap in multi-modal capabilities and artificial intelligence (AI)-human interaction. In the meanwhile, it maintains the same level of text, reasoning, and code intelligence as GPT-4-Turbo [6]. LLMs can interact with humans and the world as humans do, it is the starting point that LLMs can take over tasks from humans or collaborate naturally with humans.
Issue resolving is one of the most recent and popular focuses on solving complex tasks with LLMs. SWE-bench [1] collects $2,\!294$ real-world issues from $12$ popular Python libraries. The LLMs are tasked to resolve the issue by giving the issue description and the whole repository. This task is extremely challenging due to the need for deep reasoning about a huge amount of code and incomplete information for the task description. SWE-bench-lite [1] removes the issue with low-quality descriptions to make the task more addressable.
Since SWE-bench was released, multiple approaches have been proposed. SWE-Llama [1] adopts a pipeline with Retrieval-Augmented Generation (RAG) to generate the patch directly. Later, AutoCodeRover [7] adds code contextual retrieval with keywords in the issue description into the pipeline. It iteratively collects code context by the keywords in the issues until LLMs have collected enough information to generate a correct patch. Instead of patch generation, another approach performs iterative edits in the repository. It then uses the “git diff” command to generate patches which avoids patch format errors.
An agent-based approach for solving complex software engineering tasks for code has shown its competitiveness. One uses the multi-agent approach to automate the software development process from scratch [8]. One uses the single-agent approach to address automatic GitHub issue resolving [7, 9, 10].
To the best of our knowledge, in issue resolving scenarios, the agent-based approaches primarily focus on a single agent. Moreover, previous works perform task decomposition on-the-go, with each subsequent step being determined by the preceding one. Leveraging a multi-agent framework in issue resolving presents challenges such as: (1) Free communications between agents may lead to a non-progressing loop without termination [11]. (2) Information passed from one agent to another may incur information loss [12]. (3) Complex plans are hard to follow when multiple agents are involved. This paper aims to explore a multi-agent design called CodeR that employs a graph data structure for planning, which could mitigate the above three problems.
@@
  • Less candidate actions, easier decision. We introduce a set of diverse actions for different purposes to the agents. The number of actions is much larger compared with SWE-agent. Retaining a single-agent framework may lead to less accurate decisions. A multi-agent framework reduces the complexity of making decisions for the next action by limiting its focus to a subtask and a subset of actions.
  • Look before you leap. We believe that planning at the beginning of the pipeline is better than deciding the next steps on-the-go. Moreover, a good plan consists of small and manageable tasks that LLMs were trained to solve.
  • Bypassing instruction-following and memorization. The conventional plan generated by LLM is in the form of plain text. It is usually placed in the prompt to guide the following steps in a LLM-centered system. It requires the LLM to have a strong instruction-following ability and to have a “good” memory to execute the plan precisely and iteratively. For complex tasks, like issue resolving with complex tools, task plans in pure-text prompts will be hard to follow.
Our contributions are listed as follows:
@@
    We renew the state-of-the-art of SWE-bench lite to $28.00\%$ ($84$/$300$) with only one submission per issue.
@@ 2 Framework
As Figure  1 shows, our design contains five agents, which could collaboratively solve GitHub issues:
  • Manager: The manager is an agent who interacts with the user directly and is in charge of the whole issue-resolving task. It has two responsibilities: (1) select a plan according to the issue description. The plan specifies the agents evolved and how they should interact to finish the task. (2) Interpret the execution summary of a plan. If the execution summary has indicated that the issue has been solved. It will summarize the changes and submit a patch. If not, it will come up with a new plan or give up.
@@ 3.1 Task Graphs for Planning
This design of plans decouples agent design with the task decomposition. When designing the agents, one can only focus on the high-level goal of a sub-task without considering the details of the diverse approaches. The diversity of approaches can be specified and adjusted in the field of “task” and “downstream”. In this way, the plans can be added, deleted, and tuned without changing a single line of code for agents.
@@
We have designed four plans as Figure 3 shows. Plan A is shown in Figure 1, which is a standard flow to resolve an issue. It has no loop for simplicity and robustness. Plan B adds a loop that allows the feedback from testing. This circle is also used by Aider [14] with tests that are not related to the issue (which is also called “integration tests”). Plan C tries to resolve the issue directly for simple issues. Plan D takes a test-driven approach with a ground truth test for issues (such as “fail-to-pass” and “pass-to-pass” tests in SWE-bench). We can easily add new roles and new task plans under this framework.
@@ 4.1 Experimental Setup
SWE-bench [1] is a benchmark that can test systems’ ability to solve GitHub issues automatically. The benchmark consists of $2,\!294$ Issue-Pull Request (PR) pairs from $12$ popular open-source Python repositories (e.g., flask, numpy, and matplotlib). SWE-bench’s evaluation can be executed by providing unit test verification using post-PR behavior as the reference solution. SWE-bench lite [1] is a subset of SWE-bench, which is curated to make evaluation less costly and more accessible. SWE-bench lite comprises $300$ instances that have been sampled to be more self-contained, with a focus on evaluating functional bug fixes. More details of SWE-bench lite can be seen at https://www.swebench.com/lite.html. This paper focuses on SWE-bench lite for faster, easier, and more cost-effective evaluation.
@@
Recently, several commercial products have been released, but their technical details have not been disclosed. The following describes only their functionalities.
@@ 4.1.1 Implementation Details
In our multi-role framework, each role is considered a distinct agent with its own experimental settings, which include the model and history process window size. All roles are provided access to GPT4-preview-1106. The Manager role utilizes nucleus sampling during inference with the temperature parameter [^10] For the same input, API requests with 00 temperature may still return different content each time when the input is multi-turn. set to 00 and top_p to $0.95$. It employs full history with a file viewer’s window size of $100$. The Reproducer role similarly uses nucleus sampling, but only incorporates the last five histories. Both the Fault Localizer and Tester roles follow the same settings as the Reproducer. Finally, the Programmer role, while sharing the same nucleus sampling parameters, includes a demo in addition to the last five histories and a file viewer’s window size of $100$. This setup ensures a reduction in repetition and maintains the unique functionality of each role. In addition, we set the maximum cost to $8$\$ per issue.
@@
In fact, it is impossible to have a consistent evaluation environment for all currently proposed approaches. We make some adaptations to the evaluation environment released by AutoCodeRover [7] and use it as our evaluation environment. We reproduce all other approaches with our environment for fairness. However, the evaluation on repository “astropy” and “request” still has some environmental problems remaining. In our inference environment, commands like “edit” occasionally trigger a “container crashed” error which interrupts the process. If this occurs, we restart from the beginning of the pipeline for this issue. We pre-construct an environment-completed docker image offline to avoid wasting time on real-time installation during inference. Additionally, we divide the SWE-bench lite into six processes for parallel inference to further accelerate this process. When Fault Localizer runs the repository’s integration unit tests, it sometimes adds or modifies files within the repository, and we manually restore these files after the localization process.
@@ 4.2 Results
Table 4 displays SWE-bench lite’s results of CodeR and its comparative methods. The results show that CodeR establishes a new benchmark record on SWE-bench lite, achieving the best performance to date, compared with all other commercial products and methods. In SWE-bench lite, CodeR resolves $28.00\%$ issues at one attempt, addressing $84$ of $300$. In contrast, SWE-agent + GPT 4 and Aider solve $16.67\%$ and $24.67\%$ respectively. This proves that CodeR’s meticulously designed roles and actions are highly effective.
We notice that directly enabling LLMs to generate patches (explicit patch generation) for issues is inferior to having LLMs edit the code repository (implicit patch generation). Compared with CodeR’s $28.00\%$ resolved rate, RAG+GPT 4 and AutoCodeRover solve $2.67\%$ and $16.67\%$ respectively. We also notice that existing LLMs may struggle to generate applicable and high-quality patches, as a correct patch requires a strict format and is sensitive to line numbers, which LLMs cannot perfectly handle.
The result also shows that CodeR sends more requests, resulting in increased tokens and cost at an acceptable rate. This could be due to our fine-grained design of multi-role and actions. The $11.33\%$ improvement over SWE-agent +GPT 4 (reproduced) demonstrates that pre-planning at the beginning of the pipeline is superior to deciding the next steps on-the-go. CodeR preemptively devises multiple plans in the form of structured graphs, and all agent roles will execute the pre-defined plan strictly according to the graphs. CodeR’s leading performance also validates the effectiveness of this idea. Pre-planning also possesses a clear advantage of bypassing imperfect instruction-following and long-context memorizing abilities of LLMs. Although CodeR has achieved impressive performance, we still believe that designing a more sophisticated plan will yield more significant improvements in the future.
We also conduct ablation studies on $50$ issues of SWE-bench lite. The results in Table 5 show that removing the multi-agent would reduce CodeR’s resolved rate from $22\%$ to $10\%$. This further demonstrates that our carefully designed roles motivated by real-world company collaboration are highly useful for issue resolving tasks. Additionally, we observe a performance drop and a cost increase when we remove the fault localization action, which highlights the significant potential of combining LLMs with traditional software engineering strategies for addressing complex downstream tasks.
@@
CodeR (ours) 28.00 (84) 30.39 299K/\$3.09

### v3

@@ Abstract
GitHub issue resolving recently has attracted significant attention from academia and industry. SWE-bench [1] is proposed to measure the performance in resolving issues. In this work, we propose CodeR, which adopts a multi-agent framework and pre-defined task graphs to Repair & Resolve reported bugs and add new features within code Repository. On SWE-bench lite, CodeR is able to solve $28.33\%$ of issues, when submitting only once for each issue. We examine the performance impact of each design of CodeR and offer insights to advance this research direction [^1] https://github.com/NL2Code/CodeR.
@@ 1 Introduction
The rapidly growing capability of Large Language Models (LLMs) is dramatically reshaping many industries [2, 3, 4]. The most recent release of GPT-4o [5] demonstrates a significant leap in multi-modal capabilities and artificial intelligence (AI)-human interaction, whilst maintaining the same level of text generation, reasoning, and code intelligence as GPT-4-Turbo [6]. LLMs can interact with humans and the world as humans do, it is considered a starting point for LLMs to take over tasks from humans or collaborate naturally with humans.
Issue resolving is one of the software engineering tasks experimented with LLMs that is particularly relevant in practice. SWE-bench [1] collects $2,\!294$ real-world issues from $12$ popular Python libraries. The LLMs are tasked to resolve the issues based on the given issue description and the whole repository. This task is extremely challenging due to the need for deep reasoning about a huge amount of code and incomplete information for the task description. SWE-bench-lite [1] removes the issue with low-quality descriptions to make the task more addressable, and yet it remains highly non-trivial.
Since SWE-bench was released, multiple approaches have been proposed. SWE-Llama [1] adopt a pipeline with Retrieval-Augmented Generation (RAG) to generate the patch directly. Later, AutoCodeRover [7] added code contextual retrieval with keywords in the issue description into the pipeline. It iteratively collects code context by the keywords in the issues until LLMs have collected enough information to generate a correct patch. Instead of explicitly patch generation, SWE-agent [8] performs iterative edits in the repository. It then uses the “git diff” command to generate patches which avoids patch format errors.
In the literature on applying LLMs for solving software engineering tasks, multiple agent-based approaches have shown their competitiveness. For instance, MetaGPT [9] uses the multi-agent approach to automate the software development process from scratch. AutoCodeRover [7] and SWE-agent [8] use the single-agent approach to address automatic GitHub issue resolving.
To the best of our knowledge, in issue resolving scenarios, the agent-based approaches primarily focus on a single agent. Moreover, previous works perform task decomposition on-the-go, with each subsequent step being determined by the preceding one. Multi-agent possesses the advantage of better decoupling each role and leveraging contextual information. However, implementing a multi-agent framework in issue resolving presents challenges such as: (1) Free communications between agents may lead to a non-progressing loop without termination [10]. (2) Information passed from one agent to another may incur information loss [11]. (3) Complex plans are hard to follow when multiple agents are involved. We remark that these problems are not unlike those when human developers collaborate. In this work, we develop a multi-agent design called CodeR that effectively addresses the above mentioned problems.
@@
  • Less candidate actions, easier decision. We introduce a set of diverse actions for different purposes. The number of actions is much larger compared with the single-agent framework such as SWE-agent. To address the problem of the large number of actions, we reduce the complexity of making decisions for the next action by limiting each agent’s focus to a subtask and a subset of associated actions.
  • Look before you leap. We believe that planning at the beginning of the pipeline is better than deciding the next steps on-the-go. Moreover, a good plan should consist of small and manageable tasks that LLMs were trained to solve.
  • Bypassing instruction-following and memorization. The conventional plan generated by LLM is in the form of plain text. It is usually placed in the prompt to guide the subsequent steps in a LLM-centered system. It requires the LLM to have a strong instruction-following ability and to have a “good” memory to execute the plan precisely and iteratively. For complex tasks, like issue resolving with complex tools, task plans in pure-text prompts will be hard to follow. Therefore, we introduce a new data structure namely task graph that can ensure that all pre-designed plans are accurately followed and executed.
Our contributions are as follows:
@@
    We renew the state-of-the-art of SWE-bench lite to $28.33\%$ ($85$/$300$) with only one submission per issue.
@@ 2 Framework
As Figure  1 shows, our design contains five agents, which can collaboratively solve GitHub issues:
  • Manager: The manager is an agent who interacts with the user directly and is in charge of the whole issue-resolving task. It has two responsibilities: (1) selecting a plan according to the issue description. The plan specifies the agents evolved and how they should interact to finish the task. (2) interpreting the execution summary of a plan. If the execution summary has indicated that the issue has been solved, it will summarize the changes and submit a patch; if not, it will come up with a new plan or give up.
@@ 3.1 Task Graphs for Planning
This design of plans decouples agent design with the task decomposition. When designing the agents, one can only focus on the high-level goal of a sub-task without considering the details of the diverse approaches. The diversity of approaches can be specified and adjusted in the field of “task” and “downstream”. In this way, the plans can be easily added, deleted, and tuned without changing a single line of code for agents.
@@
We have designed four plans as Figure 3 shows. Plan A is shown in Figure 1, which is a standard flow to resolve an issue. It has no loop for simplicity and robustness. Plan B tries to resolve the issue directly for simple issues. Plan C adds a loop that allows the feedback from testing. This circle is also used by Aider [12] with tests that are not related to the issue (which is also called “integration tests”). Plan D takes a test-driven approach with a ground truth test for issues (such as “fail-to-pass” and “pass-to-pass” tests in SWE-bench). In our experiments, we use only Plan A and B for cost savings and fast evaluation.
@@ 4.1 Experimental Setup
SWE-bench [1] is a benchmark that can test systems’ ability to solve GitHub issues automatically. The benchmark consists of $2,\!294$ Issue-Pull Request (PR) pairs from $12$ popular open-source Python repositories (e.g., flask, numpy, and matplotlib). SWE-bench’s evaluation can be executed by providing unit test verification using post-PR behavior as the reference solution. SWE-bench lite [1] is a subset of SWE-bench, which is curated to make evaluation less costly and more accessible. SWE-bench lite comprises $300$ instances that have been sampled to be more self-contained, with a focus on evaluating functional bug fixes. More details of SWE-bench lite can be seen at https://www.swebench.com/lite.html. In this work, we focus on SWE-bench lite for faster, easier, and more cost-effective evaluation.
@@
Recently, several commercial products addressing issue resolving have been released, but their technical details have not been disclosed. The following describes their functionalities.
@@
  • Bytedance MarsCode Agent [^9] https://www.marscode.com is an AI coding assistant powered by GPT-4o, developed by ByteDance. Designed for multi-language support within IDE environments, it can reset repositories to undo previous modifications.
@@ 4.1.1 Implementation Details
In our multi-agent framework, each role is considered a distinct agent with its own experimental settings, which include the model and history process window size. All roles are provided access to GPT4-preview-1106. The Manager role utilizes nucleus sampling during inference with the temperature parameter set to 00 and top_p to $0.95$. It employs full history with a file viewer’s window size of $100$. The Reproducer role similarly uses nucleus sampling, but only incorporates the last five histories. Both the Fault Localizer and Tester roles follow the same settings as the Reproducer. Finally, the Programmer role, while sharing the same nucleus sampling parameters, includes a demo in addition to the last five histories and a file viewer’s window size of $100$. This setup ensures a reduction in repetition and maintains the unique functionality of each role. In addition, we set the maximum cost to $8$\$ per issue.
@@
In fact, it is impossible to have a consistent evaluation environment for all currently proposed approaches. We make some adaptations to the evaluation environment released by AutoCodeRover [7] and use it as our evaluation environment. We reproduce all other approaches with our environment for fairness. However, the evaluation on repository “astropy” and “request” still has some environmental problems remaining. In our inference environment, commands like “edit” occasionally trigger a “container crashed” error which interrupts the process. If this occurs, we restart from the beginning of the pipeline for this issue. We pre-construct an environment-completed docker image offline to avoid wasting time on real-time installation during inference. Additionally, we divide the SWE-bench lite into six processes for parallel inference to further accelerate this process. When Fault Localizer runs the repository’s integration unit tests, it sometimes adds or modifies files within the repository, and we restore these files after the localization process.
@@ 4.2 Results
Table 4 shows CodeR’s performance on SWE-bench lite and its comparative methods. The results show that CodeR establishes a new benchmark record on SWE-bench lite, achieving the best performance to date, compared with all other commercial products and methods. In SWE-bench lite, CodeR resolves $28.33\%$ issues at one attempt, addressing $84$ of $300$. In contrast, SWE-agent + GPT 4 and Aider solve $18.00\%$ and $26.33\%$ respectively. This proves that CodeR’s meticulously designed roles and actions are highly effective.
We notice that directly enabling LLMs to generate patches (explicit patch generation) for issues is less effective than having LLMs edit the code repository (implicit patch generation). While CodeR achieves $28.33\%$ resolved rate, RAG+GPT 4 and AutoCodeRover only solve $2.67\%$ and $19.00\%$ respectively. Furthermore, we observe that existing LLMs may struggle to generate applicable and high-quality patches, as a correct patch requires a strict format and is sensitive to line numbers, which LLMs cannot perfectly handle.
The result also shows that CodeR sends more requests, resulting in increased tokens and cost at an acceptable rate. This could be due to our fine-grained design of multi-role and actions. The $10.33\%$ improvement over SWE-agent +GPT 4 (reported) demonstrates that pre-planning at the beginning of the pipeline is superior to deciding the next steps on-the-go. CodeR preemptively devises multiple plans in the form of structured graphs, and all agent roles will execute the pre-defined plan strictly according to the graphs. CodeR’s leading performance also validates the effectiveness of this idea. Pre-planning also possesses a clear advantage of bypassing imperfect instruction-following and long-context memorizing abilities of LLMs. Although CodeR has achieved impressive performance, we still believe that designing a more sophisticated plan will yield more significant improvements in the future.
We also conduct ablation studies on $50$ issues of SWE-bench lite. The results in Table 5 show that removing the multi-agent & task graph would reduce CodeR’s resolved rate from $22\%$ to $10\%$. This further demonstrates that our carefully designed roles motivated by real-world company collaboration are highly useful for issue resolving tasks. Additionally, we observe a performance drop and a cost increase when we remove the fault localization action, which highlights the significant potential of combining LLMs with traditional software engineering strategies for addressing complex downstream tasks.
@@
CodeR (reported) 28.33 (85) 30.39 299K/\$3.09
CodeR (ours) 27.33 (82) 30.39 299K/\$3.09